# Personalization Toolkit: Training Free Personalization of Large Vision Language Models

### v2

@@ Abstract
Large Vision Language Models (LVLMs) have significant potential to provide personalized assistance by adapting to the unique needs and preferences of individual users. The personalization of LVLMs has emerged as a field that focuses on customizing models to recognize specific object instances and provide tailored responses. However, current methodologies depend on time-consuming test-time training for each user and object, which proves to be impractical. This paper introduces a novel, training-free approach to LVLM personalization by leveraging pre-trained vision foundation models to extract distinct features, retrieval-augmented generation (RAG) techniques to recognize instances in the visual input, and visual prompting methods. Our model-agnostic vision toolkit enables flexible and efficient personalization without the need for extensive retraining. We demonstrate state-of-the-art results, surpassing conventional training-based approaches, and set a new benchmark for LVLM personalization.
@@ 1 Introduction
Large Vision Language Models (LVLMs) [23, 22, 7, 41, 20, 1, 36] have demonstrated impressive capabilities in reasoning about visual content and answering visual questions across various domains. This suggests a great potential for deployment as visual assistants that can aid users in their daily lives. However, current LVLMs are designed to provide generic, user-independent responses and recognize objects at the category level. This limits their ability to refer to specific instances by their unique names. Figure 1 shows an example of LVLM generic inference - i.e. giving a generic response to a user query.
The task of personalizing vision-language models was recently introduced by [2] to enable LVLMs to recognize specific object instances and answer relevant questions accordingly. However, existing approaches [2, 25] require training the model for each personalized object, which can be computationally expensive and limits the model’s ability to incrementally learn new personalized concepts.
In this work, we build upon the strengths of pre-trained vision foundation models, the emerging capabilities of LLMs in in-context learning [9] and retrieval-augmented generation (RAG) [17]. Our training-free approach localizes instances using open-world object detectors [24, 16, 26] and stores reference instance-level features in memory banks, alongside their name and context. During inference, our retrieval module queries the memory bank and visual-prompts the LVLM. By storing only feature embeddings, we avoid privacy concerns derived from training on user images. An overview of our approach is shown in Figure 1.
In addition to existing benchmark datasets, we evaluate our approach using a video personalization dataset, converted to an LVLM personalization format, and further augmented with a challenging evaluation set, which reflects a realistic personalization scenario and highlights the difficulty of the personalization task. It also shows the effectiveness of patch-level features for robust instance personalization. We also discuss the limitations of existing benchmark datasets and highlight opportunities for future research.
The key contributions of our approach are: 1) We demonstrate that LVLM personalization can be achieved without training, enabling fast and efficient deployment. 2) We propose a flexible approach that leverages pre-trained vision foundation models, RAG, and visual prompting to personalize any LVLM. 3) We achieve state-of-the-art results on three personalization benchmarks, outperforming existing approaches. 4) We create a challenging evaluation set that pushes the limits of existing approaches and highlights opportunities for future research.
In the following, we discuss the related work in Section 2 and present our vision toolkit for LVLMs personalization in Section 3, evaluate our approach in Section 3 and conclude in Section 5.
@@ 2 Related Work
Personalization of Large Language Models (LLMs). LLMs personalization [40] enhances user-specific interactions without extensive fine-tuning. Approaches like prompt tuning, in-context learning, and Retrieval-Augmented Generation (RAG) offer dynamic personalization without additional model storage [21, 4, 42]. Our method avoids storing new models or adapters and is training-free.
Text-to-Image Personalization. Personalizing text-to-image generation has been explored extensively, with early methods like Textual Inversion [10], Dreambooth [31], and HyperDreamVbooth [32] requiring training for each identity, leading to scalability issues. Recent approaches [33, 39, 13] avoid test-time fine-tuning by pretraining for personalization. However, we argue that, while personalized image generation does require training, LVLMs can be personalized without changing the large language model.
Large Vision Language Models Personalization The task of personalized vision language models was introduced in MyVLM [2] by training a concept head for specific objects on top of the CLIP cls token. In a similar fashion as Dreambooth [31], MyVLM employs rare tokens to encode personalized concepts. This may have unintended consequences for language assistants, and requires optimizing the LLM caption loss for personalized conversations. Yo’LLaVA [25] adds extra tokens to the LLM head for each personalized object, learning concept tokens to describe them. Although it surpasses MyVLM in performance, the additional tokens create a challenging incremental classification problem [8]. Both models need test-time training for each new concept, limiting them to personalizing one concept at a time. Two recent works [12, 27] attempt to eliminate test-time training. While Hao et al. [12] approaches the task by large-scale pretraining on personalized conversations, PLVLM [27] pretrains an alignment module to incorporate the CLS token and image embeddings from DINOv2 [26] into the LLM. Hao et al. [12] includes a reference image in the input to the LVLM, which may limit the capability to process multiple images and hinder in-context learning. PLVLM [27] relies on querying for a specific personalized object, limiting its applicability to only VQA. Our approach applies to both VQA and captioning (general and personalized) and is the first training-free personalization solution, featuring a modular, plug-and-play design that eliminates the need for retraining or adaptation.
Vision Foundation Models. Our approach uses vision foundation models, which have become powerful generalist models through large-scale pretraining [3]. Vision-language models like CLIP [29] and Align [15] map objects to higher-level semantics with contrastive pretraining. Self-supervised models like DINO [6, 26] detect variations within the same class. Grounding DINO [24] redefines object detection as phrase grounding, and SAM [16] excels in image segmentation.
@@ 3 Approach
This section outlines our personalization toolkit, coined as PeKit, for enabling any LVLM to perform personalized object detection and answer generation. We employ a three-stage pipeline: View Extraction to extract robust object-level features from reference images and store them in a memory module, Personalized Objects Retrieval to identify objects in the query image, and Personalized Answer Generation via visual prompting. We refer to Fig 1 for an illustration of our approach.
@@ 3.2 Training-free View Extraction
Existing LVLM personalization techniques depend on image-level representations of the objects’ training views [2, 25], which can lead to overfitting to the background of each object in the reference images, particularly for training-based approaches. To avoid such a bias, our method localizes the object in the image and extracts only its corresponding features. We achieve this by utilizing an open-vocabulary segmentation network $F_{\text{ext}}$ to extract object-level masks $S_{p}$ based on each object’s generic category $k_{p}$ which can be deducted from the name or the context [^1] We refer to the Appendix for further discussion. .
@@ 3.3 Personalized Object Retrieval
where $\operatorname{sim}(E_{l},\mathbf{e}_{o_{i}})=\mathop{\mathrm{argmax}}_{k}\left\{\operatorname{sim}(\mathbf{e}_{l}^{k},\mathbf{e}_{o_{i}})\right\}$. We calculate the proposal’s similarity to the embeddings of the training views for all personalized objects. Any similarity measure (e.g., cosine similarity) can be employed for this purpose. We set a constant threshold $\tau$ to identify the personalized objects. We discard the object proposals in which no matching object is found by the retrieval module $\mathcal{R}$. Our method inherently supports the detection of multiple personalized objects (Fig 2).
@@ 3.4 Personalized Answer Generation
Once a personalized object is identified, our method generates captions specifically about that object, distinct from general captions a standard LVLM would produce. This involves emphasizing the detected object and incorporating prior knowledge about it. We achieve this through visual prompting by overlaying bounding boxes on the image and querying the LVLM to generate captions or answer questions focused on these objects. We use distinctive colors to differentiate recognized objects. We provide the LVLM with the object identifier $n_{j}$ (e.g., the instance name) and possibly a context $c_{j}$ for each personalized object. The LVLM incorporates this context and responds to queries using the given name $n_{j}$. For multiple personalized objects, we instruct the LVLM for each bounding box, name, and context. Figure 2 illustrates how the LVLM in our approach responds to the overlaid bounding boxes and our instruction. We refer to the Appendix for the exact prompt format. Note that generating answers using a context pool with varied information (e.g., monument history) is inherently supported by our method.
Figure 2: Multi-concept Personalization with Visual Prompting: PeKit handles multiple personalized objects without additional requirements. The LVLM in PeKit uses the specified bounding boxes (via visual prompts) and instance identities, with optional context (here imaginary), for personalized caption generation.
@@ 3.5 Choice of Vision Tools
We deploy GroundedSAM [30] as the open-vocabulary segmentation network $F_{\text{ext}}$ and ablate using GroundingDINO [24], where the mask is represented by the object’s bounding box. We use DINOv2 [26] as the image encoder $F_{\text{emb}}$ to extract patch-level features of the personalized objects. Image encoders trained with a self-supervised objective, such as DINOv2, produce distinctive features that improve the re-identification of personalized objects as we empirically show in the Appendix in a comparison with CLIP [28]. During inference, GroundingDINO [24] is queried with the generic term ‘object’ to generate object proposals.
3.6 Summary
To summarize our approach, there are two phases
Personalized object introduction: The user introduces an object to the LVLM with one or more reference images, name, and optional context. We extract robust patch-level features and store the object in our memory module $\mathcal{M}$ as $(E_{i},(n_{i},c_{i}))$.
Personalized Inference: Given an image and a question $Q$, we extract object proposals and query our retrieval module $\mathcal{R}$ with the bounding boxes features. If a match or multiple matches are found $\{(E_{i},(n_{i},c_{i}))\}$, we highlight the bounding box(es) on the query image with unique color(s) and instruct the LVLM to respond using $\{(n_{j},c_{j})\}$. If no match is found, the LVLM responds without further instructions.
@@ 4 Experiments
Figure 3: Our new evaluation set This-Is-My-Img based on This-Is-My dataset [38] : Training and validation sets examples for the category Reynard’s Work Chair.
This section evaluates our approach against existing methods for vision language personalization.
Datasets. MyVLM [2] dataset consists of 29 object categories and Yo’LLaVA [25] dataset includes 40 categories of objects, buildings, and people. It also features a VQA benchmark with multiple-choice questions (A/B). For both MyVLM and Yo’LLaVA datasets, we use only the reference images, discarding the ground-truth captions and negative images. This-Is-My-Img Dataset Yo’LLaVA and MyVLM datasets are simple and object-centric, lacking real-world complexity. To address this, we propose a new evaluation set for LVLM personalization based on the This-Is-My dataset [38], originally created for video-level detection of personalized instances reflecting real-world scenarios where objects may be partially visible or in the background. To make this dataset compatible with image-level training/evaluation, we process it as follows: We extract five reference frames per object from the training segments, ensuring the entire object is visible. For validation, we sample every 10th frame  [^2] Due to privacy concerns, some original video segments are unavailable, limiting our benchmark to 14 categories. See the Appendix for more details.. The validation frames are categorized into positives and negatives for each personalized object. Positives: frames containing the personalized object. Negatives: frames from the same validation segments where the object is not visible (denoted as Hard negatives) and frames from videos of other objects (Other). We further introduce an additional negative validation set (Fake) by generating images similar to the personalized objects frames with varied perspectives and environments using ChatGPT-4o (see Figure 3 for examples). We ensure that objects in the fake images differ in characteristics from the personalized objects. Our evaluation set challenges LVLM personalization methods with images containing multiple foreground objects, background elements, motion blur, and synthetic images [^3] This-Is-My-Img will be made publicly available.. We further extend This-Is-My-Img with a VQA evaluation set containing 70 multiple-choice (A/B) questions.
Implementation Details. Our method is generic regarding the choice of the LVLM model. We use LLaVA1.5-13B [22] as our primary LVLM model for consistency with previous works. We utilize Cosine Similarity with a constant threshold of $\tau$=0.75 for detecting personalized objects across all datasets (3.3). Refer to the Appendix for further details and results with automatic object-specific threshold using reference image similarity. We also report the compute and memory overhead of our personalization toolkit and compare them to the backbone in the Appendix.
@@
Evaluation Metrics. We follow the metrics and notation of Yo’LLaVA [25] and MyVLM [2], reporting Positive Accuracy (Recall), Negative Accuracy (Specificity) and Weighted Accuracy for $n$ personalized objects. We further report Precision, defined as the ratio of true positive predictions to the total number of instances predicted as positive, averaged over $n$ personalized objects. We believe this is a more comprehensive evaluation recognizing the trade-off between the aforementioned metrics and precision. We report the accuracy (percentage of correct answers) for the VQA task. Following MyVLM, we report CLIPScore (measuring the similarity between the caption and the image) and personalization recall (assessing the LVLM’s ability to incorporate the personalized object’s name into the caption) on the MyVLM dataset.
@@
This-Is-My-Img Dataset
@@ 4.1 Visual-Recognition
Table 1 presents the performance of our approach compared to current methods, MyVLM [2] and Yo’LLaVA [25], on their respective datasets. On the MyVLM dataset, our method PeKit achieves SOTA results, improving both positive and negative accuracy, with an average improvement of 1.9%. On the Yo’LLaVA dataset, our method significantly improves negative accuracy by 8.9% while having a lower positive accuracy than Yo’LLaVA [25], resulting in an average improvement of 2.5% on weighted accuracy, hence achieving best results on average. Our PeKit might miss a few challenging appearances of personalized objects but has a very low false positive rate, meaning we avoid incorrect recognition of personalized objects. We believe it is better to resort to generic captions than to provide incorrect personalized answers.
@@
This-Is-My-Img Dataset
@@
Table 3: Visual Question Answering Performance. PeKit achieves SOTA performance on the three datasets.
Next, we evaluate the performance on our introduced benchmark This-Is-My-Img in Table 2. MyVLM tends to provide positive responses to test images, resulting in very high positive accuracy but much lower negative accuracy ($14.6\%$ on other and $4.7\%$ on hard negatives). This indicates that MyVLM tends to capture the scene in which the object occurs rather than the specific object. Note that we followed exactly the training pipeline and requirements for positive and negative images of MyVLM. Yo’LLaVA has similar but less pronounced trends with high positive accuracy ($87.1\%$) and a moderate negative accuracy of $84.8\%$ on other negatives and $61.9\%$ on hard negatives. While PeKit achieves a lower positive accuracy, we demonstrate a more robust negative accuracy ($99.9\%$ and $96.0\%$ on all and hard negatives respectively). We achieve a precision of $90.1\%$ which is $48\%$ higher than Yo’LLaVA indicating that our approach refrains from providing wrong personalization. On the fake images, smaller differences appear between methods. The fake images sometimes contain very similar objects to the personalized object in terms of style, resulting in confusion by the methods, indicating room for future work. Yo’LLaVA and MyVLM detect the domain shift on some images more than PeKit. While PeKit is quite robust to different objects, it can be challenged on very similar fake and small (see limitations in Appendix). Overall, PeKit shows striking robustness and balanced behavior on both positive and negative accuracy achieving an average visual recognition performance of $82.86$ compared to 67.38 by Yo’LLaVA and 33.92 by MyVLM. Overall, this evaluation set illustrates the difficulty of the personalization task beyond object-centric images, moving closer to realistic scenarios for intelligent visual assistants. This opens the door for future research on this challenging benchmark.

### v3

@@ Abstract
Personalization of Large Vision-Language Models (LVLMs) involves customizing models to recognize specific users and object instances, and to generate contextually tailored responses. Existing approaches typically rely on time-consuming test-time training for each user or object, making them impractical for real-world deployment, a limitation reflected in current personalization benchmarks, which are focused on object-centric, single-concept evaluations. In this paper, we present a novel training-free approach to LVLM personalization and introduce a comprehensive real-world benchmark designed to rigorously evaluate various aspects of the personalization task. Our method leverages pre-trained vision foundation models to extract distinctive features, applies retrieval-augmented generation (RAG) techniques to identify instances within visual inputs, and employs visual prompting strategies to guide model outputs. Our model-agnostic vision toolkit enables efficient and flexible multi-concept personalization across both images and videos, without any additional training. We achieve state-of-the-art results, surpassing existing training-based methods.
@@ 1 Introduction
Large Vision Language Models (LVLMs) [17, 16, 4, 31, 15, 1, 28] have demonstrated impressive capabilities in reasoning about visual content and answering visual questions across various domains. This suggests a great potential for deployment as visual assistants that can aid users in their daily lives. However, current LVLMs are designed to provide generic, user-independent responses and recognize objects at the category level (Fig. 1, left).
The task of personalizing vision-language models was introduced by [2] to enable LVLMs to recognize specific object instances and answer relevant questions accordingly. Existing approaches [2, 20] rely on training for a specific personalized object, diverting the LVLM from its original capabilities and incurring a large compute cost.
In this work, we introduce a training-free approach that builds upon the strengths of pre-trained vision foundation models, the emerging capabilities of LLMs in in-context learning [7] and retrieval-augmented generation (RAG) [12]. Our method localizes instances using open-world object detectors [18, 11, 21] and stores reference instance-level features in memory banks, alongside their name and context. During inference, our retrieval module queries the memory bank and visual-prompts the LVLM. An overview of our approach is shown in Fig. 1.
Furthermore, existing personalization benchmarks primarily focus on object-centric, single-concept tasks, falling short of capturing the complexity of real-world applications. To address this gap, we introduce a novel and challenging benchmark derived from a video personalization dataset. It reflects the real-world complexity of the personalization task, with subsets targeting single-concept, multi-concept, image-based, and video-based VQA. We use this benchmark to rigorously evaluate our method across diverse scenarios.
The key contributions of our approach are: 1) We show that LVLM personalization is possible without training, enabling fast deployment. 2) We present a flexible method that supports multi-concept and video personalization across LVLMs using vision foundation models, RAG, and visual prompting. 3) We introduce a challenging benchmark that exposes current limitations and guides future research [^1] Our benchmark can be previewed in the supplementary materials and will be made publicly available.. 4) We achieve state-of-the-art performance on various tasks using both existing benchmark datasets and our proposed benchmark, consistently surpassing previous methods.
We discuss the related work in Section 2 followed by an introduction to our vision toolkit for LVLMs personalization in Section 3. We then present our real-world benchmark and evaluate our approach in Section 2, and conclude in Section 5.
@@ 2 Related Work
Large Vision Language Models Personalization The task of personalized vision language models was introduced in MyVLM [2] by training a concept head for specific objects on top of the CLIP cls token. In a similar fashion as Dreambooth [25], MyVLM employs rare tokens to encode personalized concepts. This may have unintended consequences for language assistants, and requires optimizing the LLM caption loss for personalized conversations. Yo’LLaVA [20] adds extra tokens to the LLM head for each personalized object, learning concept tokens to describe them. Although it surpasses MyVLM in performance, the additional tokens create a challenging incremental classification problem [6] when introducing new personalized concepts. Both models need test-time training for each new concept, limiting them to personalizing one concept at a time. Two recent works [9, 22] attempt to eliminate test-time training. While Hao et al. [9] approaches the task by large-scale pretraining on personalized conversations, PLVLM [22] pretrains an alignment module to incorporate the CLS token and image embeddings from DINOv2 [21] into the LLM. Hao et al. [9] includes a reference image in the input to the LVLM, which may limit the capability to process multiple images and hinder in-context learning. PLVLM [22] relies on querying for a specific personalized object, limiting its applicability to only VQA. Current approaches are mostly designed for single concept/image personalization and it is unclear how to scale such approaches to more realistic setting. To the contrary, our approach offers a training-free personalization solution, featuring a modular, plug-and-play design that eliminates the need for retraining or adaptation while naturally and efficiently scaling to multi concept and video personalization.
@@ 3 Approach
This section outlines our personalization toolkit, coined as PeKit, for enabling any LVLM to perform personalized object detection and answer generation. We employ a three-stage pipeline: View Extraction to extract robust object-level features from reference images and store them in a memory module, Personalized Objects Retrieval to identify objects in the query image, and Personalized Answer Generation via visual prompting. We refer to Fig. 1 for an illustration of our approach.
@@ 3.2 Training-free View Extraction
Existing LVLM personalization techniques depend on image-level representations of the objects’ training views [2, 20], which can lead to overfitting to the background of each object in the reference images, particularly for training-based approaches. To avoid such a bias, our method localizes the object in the image and extracts only its corresponding features. We achieve this by utilizing an open-vocabulary segmentation network $F_{\text{ext}}$ to extract object-level masks $S_{p}$ based on each object’s generic category $k_{p}$ which can be deducted from the name or the context [^2] We refer to Appendix A.5 for further discussion..
@@ 3.3 Personalized Object Retrieval
where $\operatorname{sim}(E_{l},\mathbf{e}_{o_{i}})=\mathop{\mathrm{argmax}}_{k}\left\{\operatorname{sim}(\mathbf{e}_{l}^{k},\mathbf{e}_{o_{i}})\right\}$. We calculate the proposal’s similarity to the embeddings of the training views for all personalized objects. Any similarity measure (e.g., cosine similarity) can be employed for this purpose. We set a constant threshold $\tau$ to identify the personalized objects. We discard the object proposals in which no matching object is found by the retrieval module $\mathcal{R}$. Our method inherently supports the detection of multiple personalized objects
@@ 3.4 Personalized Answer Generation
Once a personalized object is identified, our method generates captions specifically about that object, distinct from general captions a standard LVLM would produce. This involves emphasizing the detected object and incorporating prior knowledge about it. We achieve this through visual prompting by overlaying bounding boxes on the image and querying the LVLM to generate captions or answer questions focused on these objects. We use distinctive colors to differentiate recognized objects. We provide the LVLM with the object identifier $n_{j}$ (e.g., the instance name) and possibly a context $c_{j}$ for each personalized object. The LVLM incorporates this context and responds to queries using the given name $n_{j}$. For multiple personalized objects, we instruct the LVLM for each bounding box, name, and context. We refer to Appendix C.1 for the exact prompt format.
3.5 Video-personalization
To conduct personalized video question answering (VQA), we apply the previously described method to a set of sampled video frames. Each personalized object is consistently highlighted using the same bounding box color across all frames to maintain identity. These sampled frames—including those with visual prompts (bounding boxes) and those without any detected personalized objects—are then fed into a video large language model.
The model is prompted to generate captions or answer questions using each detected object’s identifier $n_{j}$, an optional context $c_{j}$, and the unique bounding box color used for the visual prompting of that object instance (Appendix C.1).
@@ 3.6 Choice of Vision Tools
We deploy GroundedSAM [24] as the open-vocabulary segmentation network $F_{\text{ext}}$ and ablate using GroundingDINO [18], where the mask is represented by the object’s bounding box. We use DINOv2 [21] as the image encoder $F_{\text{emb}}$ to extract patch-level features of the personalized objects. Image encoders trained with a self-supervised objective, such as DINOv2, produce distinctive features that improve the re-identification of personalized objects as we empirically show in Appendix A.1 in a comparison with CLIP [23]. During inference, GroundingDINO [18] is queried with the generic term ‘object’ to generate object proposals.
@@ 4 Experiments
Figure 2: Our proposed evaluation set This-Is-My-Img, built on the This-Is-My dataset [30]: Example reference views and validation samples from the single-concept category Reynard’s Work Chair and the multi-concept category Nikki-Nikki’s Car. Faces are blurred to ensure compliance with GDPR.
We present our LVLM personalization validation benchmark and compare our approach with existing personalization methods.
4.1 Real-world Personalization Benchmark
Prior personalization datasets such as MyVLM[2] and Yo’LLaVA[20] primarily focus on object-centric scenarios in controlled settings. MyVLM [2] dataset consists of 29 object categories and Yo’LLaVA [20] dataset includes 40 categories of objects, buildings, and people. It also features a VQA benchmark with multiple-choice questions (A/B). Despite their utility, these datasets lack real-world complexity—such as background clutter, partial occlusions, and diverse contexts—limiting their effectiveness for evaluating robust LVLM personalization.
To address this gap, we introduce a more realistic evaluation set called This-Is-My-Img benchmark, based on the This-Is-My dataset [30], originally developed for video-level detection of personalized instances in realistic settings where objects may be partially occluded, in the background, or poorly lit. We design several splits to evaluate specific capabilities of personalization methods.
Reference Views: As in prior work, we extract five reference images per object from training segments. Unlike earlier datasets, objects in these frames may appear under suboptimal conditions—e.g., partially visible, in the background, or poorly lit (Fig. 3 Appendix).
Single-Concept Eval Set: Set of images with a single personalized object. Fig. 2 shows examples. For each personalized object, images are labeled as:
  • Positive: Contain the personalized object and are used to evaluate the method’s recall performance.
  • Negative-Hard: Taken from validation segments of the same object where the object is not visible. This tests methods’s overfitting to the background associated with the personalized object.
  • Negative-Other: Sourced from validation segments of other personalized objects. Measures method’s robustness to intra-instance and the general dataset bias.
  • Negative-Fake: AI-generated images (via GPT) that mimic the personalized object under varied perspectives and environments. These samples are used to assess the method’s robustness against high visual similarity combined with a distribution shift.
  • Single-Concept VQA: A set of A/B multiple-choice questions based on positive images, in Yo’LLaVA-style format.
Multi-Concept Set: Images including a pair of personalized concepts. For example a person-object pair (Nikki & Nikki’s car) or a person-person pair (Casey & Casey’s son). To achieve this, we extend the personalized instances of the original dataset. Positives of a pair of personalized objects are images with both instances of the pair appearing together while negatives include all images from other pairs in the dataset. See Fig. 2 for examples.
  • Multi-Concept VQA: Provides open-ended questions for each image in the multi-concept set that address both objects depicted.
Video-QA Set: Original video clips from the This-Is-My dataset are paired with open-ended questions about the personalized object in the video, testing both temporal reasoning and the method’s adaptability to video inputs.
Our evaluation benchmark presents a comprehensive challenge for LVLM personalization, requiring recognition and reasoning over personalized objects in realistic, cluttered, and dynamic conditions. Appendix.B presents more details on each set.
4.2 Metrics and Evaluation Setting
@@
Single-Concept Metrics. We follow the metrics and notation of Yo’LLaVA [20] and MyVLM [2], reporting Positive Accuracy (Recall), Negative Accuracy (Specificity) and Weighted Accuracy for $n$ personalized objects. We further report Precision, defined as the ratio of true positive predictions to the total number of instances predicted as positive, averaged over $n$ personalized objects. We believe this is a more comprehensive evaluation recognizing the trade-off between the aforementioned metrics and precision. VQA accuracy on single-concept datasets is measured as the percentage of correctly selected multiple-choice questions. In line with MyVLM, we also report CLIPScore, which evaluates the similarity between the generated caption and the image, and Personalization Recall, which measures how effectively the model incorporates the personalized object’s name into the caption on the MyVLM dataset.
Multi-Concept Metrics. Positive accuracy is measured as to number of times when both personalized objects in the image are correctly detected. Negative accuracy is measured as the number of times at least one of the two objects are correctly not detected in an image.
Video and Multi-concept VQA Metric. For open-ended VQA validation, we follow a protocol similar to that of [19], utilizing ChatGPT-3.5 to evaluate the accuracy of the model’s responses (see Appendix C.2 for more details).
Implementation Details. The modularity of our method makes it generic regarding the choice of the LVLM model. We use LLaVA1.5-13B [16] as our primary LVLM model for consistency with previous works. We pair PeKit with InternVL2-26B [4] , a state-of-the-art LVLM to ablate. We use LLaVA-OneVision-Qwen2-7B as our default video model and ablate with InternVideoChat2.5 as a superior video model. We utilize Cosine Similarity with a constant threshold of $\tau$=0.75 for detecting personalized objects across all datasets (3.3). Refer to Appendix A.7 for further details on compute and memory overhead of our personalization toolkit.
@@
This-Is-My-Img Single-Concept Dataset
@@
This-Is-My-Img Multi-Concept Dataset
Method/Metric Precision Accuracy Avg.
Positive Negative
Yo’LLaVA 10.0 83.6 25.0 54.3
PeKit 96.1 45.4 99.8 72.6
@@ 4.3 Visual-Recognition
Previous Datasets Tab. 1 compares our method to MyVLM [2] and Yo’LLaVA [20] on their respective benchmarks. On MyVLM, PeKit achieves a new SOTA, improving both positive and negative accuracy with an average gain of 1.9%. On Yo’LLaVA, it increases negative accuracy by 8.9%, with slightly lower positive accuracy, resulting in a SOTA 2.5% improvement in weighted accuracy. While PeKit may miss some difficult instances, its low false positive rate helps avoid incorrect personalized outputs, favoring generic captions when uncertain. We ablate the choice of an open-world object detector, G-DINO [18], compared to the open-world semantic segmentation model, G-SAM [24], for $F_{\text{ext}}$ in Eq. 1. The results show that our method outperforms previous methods in both cases. However, the more precise segmentation model, which extracts only patches of the object of interest, achieves the best accuracy on average.
@@
This-Is-My-Img Single-Concept Dataset
@@
This-Is-My-Img Multi-Concept Dataset
LLaVA 49.0
Yo’LLaVA 12.7
PeKit (LLaVA) 56.3
PeKit (InternVL) 63.6
This-Is-My-Video Dataset
LLaVA-OneVision-Qwen2 23.0
InternVideoChat2.5 56.6
PeKit (LLaVA-OneVision-Qwen2) 35.0
PeKit (InternVideoChat2.5) 61.3
Table 3: Visual Question Answering Performance. PeKit achieves SOTA performance on all benchmarks.
Figure 3: Qualitative Results: PeKit handles a range of personalization tasks, encompassing both single- and multi-concept personalization in images and videos. For video personalization, the VLM model can reliably track the target object across frames using only a few confidently annotated instances. One representative frame is shown per scene. Faces are blurred to ensure compliance with GDPR.
This-Is-My-Img Benchmark. Tab. 2 reports the performance on our benchmark. For single-concept images, MyVLM shows high positive accuracy but poor negative accuracy (14.6% on other and 4.7% on hard negatives), suggesting it focuses on scene context over specific objects and tends to provide mostly positive responses to test images. Yo’LLaVA performs similarly but better, with 87.1% positive accuracy and 84.8% on other and 61.9% on hard negative samples. In contrast, PeKit offers stronger negative accuracy (99.9% and 96.0%) with a precision of 90.1%—48% higher than Yo’LLaVA—indicating fewer false positives.
On fake images, differences narrow. All models struggle with objects stylistically close to the personalized one with image-level methods like MyVLM and Yo’LLaVA potentially having a higher chance to detect the domain shift. However, PeKit also demonstrates robustness to domain variations, despite some difficulty with highly similar objects. Overall, PeKit achieves the best balance, with an average accuracy of 82.86%, outperforming Yo’LLaVA (67.38) and MyVLM (33.92).
In multi-concept settings, PeKit shows slightly lower accuracy in dual-object detection but achieves much higher precision and negative accuracy, leading to an 18.3% overall accuracy gain over Yo’LLaVA. These results highlight our training-free method’s superior robustness and lower false positive rate, especially in complex real-world scenarios.
Overall, this evaluation set illustrates the difficulty of the personalization task beyond object-centric images, moving closer to realistic scenarios for intelligent visual assistants. This opens the door for future research on this challenging benchmark.